How many grams of NH3 will be produced, if 97.77g of CaCO3 are also produced in the reaction below? Show the solution

1
Expert's answer
2022-05-31T20:34:05-0400

If the reaction is CaCN2 + 2H2O = CaCO3 + 2NH3 then

n(CaCO3) = m/M

n(CaCO3) = 97.77/100 = 0.9777 mole

n(NH3) = 1.955 mole

m(NH3) = 1.955•17 = 33.24 g